OpenStudy (anonymous):

You go to the doctor and he gives you 13 milligrams of radioactive dye. After 24 minutes, 4.25 milligrams of dye remain in your system. To leave the doctor's office, you must pass through a radiation detector without sounding the alarm. If the detector will sound the alarm if more than 2 milligrams of the dye are in your system, how long will your visit to the doctor take, assuming you were given the dye as soon as you arrived? Give your answer to the nearest minute. You will spend ? minutes at the doctor's office.

OpenStudy (anonymous):

56 min.

OpenStudy (anonymous):

how? did you get it?

OpenStudy (anonymous):

please:)

OpenStudy (amistre64):

pert might help

OpenStudy (anonymous):

ok:)

OpenStudy (amistre64):

4.25 = 13(e^24t)

OpenStudy (amistre64):

4.25/13 = e^24t ln(4.25/13) = 24t ln(4.25/13) ---------- = t 24

OpenStudy (amistre64):

umm... t = rate lol

OpenStudy (amistre64):

we use the rate to determine when 2 is the answer

OpenStudy (amistre64):

2 = 13(e^rate(time))

OpenStudy (amistre64):

time then equals: ln(2/13) ------- = minimum time in office rate

OpenStudy (anonymous):

ok i see...:) hopefully i can get the next story prob right:) i have 17 of them:(

OpenStudy (anonymous):

someone told me about this!:O! thanks!!

OpenStudy (anonymous):

the website i mean haha

OpenStudy (amistre64):

yeah, i just stumbled upon it one day while cleaning behind the fridge lol

OpenStudy (anonymous):

lol